# Diffusion-based Speech Enhancement with Schrödinger Bridge and Symmetric Noise Schedule

### v1

@@ IV-A Experimental Setup
SBSE $\underline{1.42_{\pm 0.09}}$ $\underline{1.63_{\pm 0.11}}$ $2.86_{\pm 0.07}$ $\underline{7.88_{\pm 0.90}}$ $\underline{11.75_{\pm 0.80}}$ $\underline{22.89_{\pm 0.52}}$ $\underline{3.78_{\pm 0.06}}$ $\underline{3.85_{\pm 0.05}}$ $\mathbf{3.93_{\pm 0.02}}$
@@
For the data preprocessing, a window size of 510 ms, a hop length of 128 ms, and the Hann window are used to transform the waveform into a complex spectrogram. We randomly select the segment that lasts 256 frames from the complex spectrogram at each training step. Following previous work [5], we apply the same amplitude transformation technique on the complex spectrogram to bring out the frequency bins with low energy, thereby balancing the data.
@@ IV-B Baselines
In evaluating our proposed speech enhancement method, we compare it with two diffusion-based models, CDiffuSE [4] and SGMSE+ [5], and two discriminative models, DeepFilterNetV3 [34] and MetricGAN+ [35]. For SGMSE+, we retrained the model with our dataset from the 2023 DNS challenge, whereas for the other baselines, we utilized the pre-trained models provided by the authors. We could not include NVIDIA SB-based baseline [11] as it was published shortly before this submission. To evaluate the influence of ratio mask in our model, we use SBSE to denote the SB model without mask input and use SBSE-M to represent the model with ratio mask input. We compared their performance.
@@ IV-C Speech Quality Assessment
  • Compared with another diffusion-based model, SGMSE+, both SBSE and SBSE-M outperform SGMSE+ in low SNR environments while achieving comparable results in high SNR scenarios. Qualitative assessments indicate that SGMSE+ produces vocalizing artifacts, such as sounds resembling breathing and sighing, in highly noisy scenarios, leading to speech distortion. Unlike SGMSE+, our approach rarely produces strong, pronounced, distorted artificial noises, thus achieving superior performance.
  • Proposed SBSE-M and SBSE models outperform the discriminative approaches across all scenarios. Notably, in low SNR conditions, our methods exceed the DeepFilterNetV3, which ranks highest among the baseline models. In high SNR environments, our models excel, producing high-quality enhanced speech and demonstrating substantial improvements over discriminative methods.
  • As shown in Fig. 2, the proposed SBSE system received the highest scores in the listening test under low SNR situations. Especially in the most challenging condition ($SNR=-5$), SBSE produced fair-quality speech, while the other methods received poor scores. These results align with those presented in Table I, indicating that our SBSE model can generate speech with better intelligence and outshines all other models in scenarios with low SNR.
  • In most scenarios, incorporating a ratio mask improves the quality of generated speech. The mask input boosts speech quality in low SNR situations by providing extra information to the network, thereby mitigating the effects of strong noise and the lack of detail in the original input. However, in high SNR environments where the original input is relatively clean, the mask’s impact is less pronounced.
@@
TABLE II: Inference speed evaluation of ten 10-second audio files measured on an NVIDIA GeForce RTX 4070 Ti.
Methods DeepFilterNet SBSE SBSE-M CDiffuSE SGMSE+
NFE 1 5 5 6 30
RTF ($\downarrow$) 0.0255 0.210 0.212 1.131 2.104
@@ IV-D Inference Speed Evaluation
Tab. II evaluates the sampling speed of our proposed methods and baseline models. The Number of Function Evaluations (NFE) for baseline models are adopted from their original paper. For our methods, we have configured the NFE to 5 for SBSE and SBSE-M. This configuration has been determined to provide satisfactory outcomes in our experiments while being computationally efficient. We use the real-time factor (RTF) to represent the inference speed, which indicates the ratio of the time to process the audio to the audio length. Compared to discriminative models, which require only one step for inference, our model trades off longer inference time for improved outcomes. Unlike diffusion-based models, SBSE operates with fewer steps, faster processing, and better qualitative results.
@@ V Conclusions
In this paper, we have introduced the Schrödinger Bridge-Based Speech Enhancement method and proposed the two-stage system integrating ratio mask information into the generative model. Our experiment results have shown that the SBSE model outperforms both discriminative and diffusion-based baseline models in low SNR conditions, achieving exceptional results in high SNR scenarios as well. Furthermore, we have demonstrated the significance of the ratio mask in enhancing speech quality under very noisy conditions. Additionally, our method outperforms other diffusion-based models in audio quality while ensuring reasonable inference time.
Although our proposed method achieves outstanding performance, it has limitations. The generative SB model occasionally produces phonetically accurate vocalizing sounds lacking linguistic meaning in extremely noisy regions where the original speech input is strongly corrupted or completely lost. The current methods fail to restore the audio fully. We suspect this hallucination issue may be inherent to the nature of generative models that should be addressed in the future.

### v2

@@ IV-A Experimental Setup
StoRM $\underline{1.43_{\pm 0.08}}$ $\underline{1.64_{\pm 0.11}}$ ${2.61_{\pm 0.07}}$ $4.45_{\pm 0.19}$ $8.84_{\pm 1.10}$ $21.50_{\pm 0.61}$ $3.32_{\pm 0.06}$ $3.42_{\pm 0.06}$ $3.66_{\pm 0.02}$
SBSE $1.42_{\pm 0.09}$ ${1.63_{\pm 0.11}}$ $2.86_{\pm 0.07}$ $\underline{7.88_{\pm 0.90}}$ $\underline{11.75_{\pm 0.80}}$ $\underline{22.89_{\pm 0.52}}$ $\underline{3.78_{\pm 0.06}}$ $\underline{3.85_{\pm 0.05}}$ $\mathbf{3.93_{\pm 0.02}}$
@@
For the data preprocessing, a window size of 32 ms, a hop length of 8 ms, and the Hann window are used to transform the waveform into a complex spectrogram. We randomly select the segment that lasts 256 frames from the complex spectrogram at each training step. Following previous work [5], we apply the same amplitude transformation technique on the complex spectrogram to bring out the frequency bins with low energy, thereby balancing the data.
@@ IV-B Baselines
In evaluating the proposed methods, the SBSE and its ratio mask extension SBSE-M models, we compare them with two discriminative models DeepFilterNetV3 [34] and MetricGAN+ [35], and three diffusion-based models, CDiffuSE [4], SGMSE+ [5] and two-stage model StoRM [12].
We did not include NVIDIA SB-based baseline [11] as it was published shortly before this submission and without pre-trained models.
@@ IV-C Speech Quality Assessment
  • Compared with StoRM and SGMSE+ diffusion-based models, the SBSE-M outperforms them in low SNR environments while achieving comparable results in high SNR scenarios. Qualitative assessments indicate that SGMSE+ produces vocalizing artifacts, such as sounds in highly noisy scenarios resembling breathing and sighing. Unlike baselines, our approach rarely produces strong, pronounced, distorted artificial noises.
  • Proposed SBSE-M and SBSE models outperform the discriminative approaches across all scenarios. Notably, in low SNR conditions, our methods exceed the DeepFilterNetV2, which ranks highest among the baseline models. In high SNR environments, our models excel, producing high-quality enhanced speech and demonstrating substantial improvements over discriminative methods.
  • As shown in Fig. 2, the proposed SBSE system received the highest scores in the listening test under low SNR situations. Especially in the most challenging condition ($SNR=-5$), SBSE produced fair-quality speech, while the other methods received poor scores. These results align with those presented in Table I.
  • In most scenarios, incorporating a ratio mask improves the quality of generated speech. The mask input boosts speech quality in low SNR situations by providing extra information to the network, thereby mitigating the effects of strong noise and the lack of detail in the original input.
@@ IV-D Inference Speed Evaluation
We also evaluated the sampling speed of our proposed methods and baseline models of ten 10-second audio files measured on an NVIDIA GeForce RTX 4070 Ti. The Number of Function Evaluations (NFE) for baseline models are adopted from their original paper. For our methods, we have configured the NFE to 5 for SBSE and SBSE-M. This configuration has been determined to provide satisfactory outcomes in our experiments while being computationally efficient. We use the real-time factor (RTF) to represent the inference speed, which indicates the ratio of the time to process the audio to the audio length. The fastest model was the discriminative DeepFilterNet with $0.026$ RTF. Our proposed SBSE and SBSE-M models were about 10 times slower with $0.21$ RTF. The generative baselines CDiffuSE and SGMSE+ achieved RTF $1.31$ and $2.11$, respectively. Compared to discriminative models, which require only one step for inference, our model trades off longer inference time for improved outcomes. Unlike diffusion-based models, SBSE operates with fewer steps, faster processing, and better qualitative results.
@@ V Conclusions
In this paper, we have revisited the Schrödinger Bridge-Based Speech Enhancement method and proposed the two-stage system integrating ratio mask information into the generative model. Our experiment results have shown that the SBSE model outperforms both discriminative and diffusion-based baseline models in low SNR conditions and not degrade signals in high SNR scenarios. Furthermore, we have demonstrated the significance of the ratio mask in enhancing speech quality under very noisy conditions. Additionally, our method is also faster compared to other diffusion-based models.
Although our proposed method achieves promising performance, it has limitations. The generative SB model occasionally produces phonetically accurate vocalizing sounds lacking linguistic meaning in extremely noisy regions; the current methods fail to restore the audio fully, which belongs to our future work.